OpenStudy (anonymous):

What symbol should be placed in the blank to make the equation true? 3 + 10___ 2 - 4 = 4 A)addition sign B)subtraction sign C)multiplication sign D)division sign

OpenStudy (anonymous):

@Hope_nicole @stamp

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

D. 3+10/2-4=4 3+5-4=4 8-4=4 4=4

OpenStudy (anonymous):

i was thinkn the same thing bcuz it cudnt b nothing else

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

If a is a negative number and b is a negative number, which expression is always negative? A) a(b) B) a ÷ b C) a - b D) a + b

OpenStudy (anonymous):

d, adding two negative will never make a positive.